The limit: limit x→0^+ (xcosecx)^x does not exist

True or false with full explanation



1
Expert's answer
2021-04-13T13:37:09-0400

Let us show that the limit "\lim\limits_{ x\to 0^+} (x\cosec x)^x" exists.

Since "\lim\limits_{ x\to 0} \frac{\sin x}{x}=1," we conclude that

"\lim\limits_{ x\to 0^+} (x\cosec x)^x=\lim\limits_{ x\to 0^+} (\frac{x}{\sin x})^x=
\lim\limits_{ x\to 0^+} (\frac{\sin x}{ x})^{-x}=1^0=1"


Answer: false